40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9103     Entry time: Wed Sep 4 17:22:09 2013
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Demod phases for RFPDs 

I checked the demod phases for AS55, POP110, and all the REFL PDs. 

AS55:  I locked MICH, and shook the ITMs (+1 for Y, -1 for X), and watched the AS55 I & Q spectra at 580Hz (notch in the servo was enabled).  I rotated the phase from -32.0 to +13.0 to get the signal entirely in the Q phase.

POP110:  I locked the PRMI (triggering on POP22), and maximized POP22.  I then rotated the phase of POP110 until the signal was maximally positive.  I forgot what the starting phase was, but it is now 84.  The POP11_I signal was entirely negative when I started, so the new phase is about 180 from the old phase.  I also checked by unlocking the cavity, and seeing that a large peak in POPDC corresponded to large negative dips in POP110_I and POP22_I.

REFL PDs:  I locked the PRMI, and shook the PRM (notches in the servos were enabled for both MICH and PRCL).  Maximized the peak in the I phase.  REFL11 was fine, REFL33 was fine.  REFL55 was changed from 120 to 45.  REFL165 was changed from 106 to 96.

I restored the SRM on the IFO_ALIGN screen, but the saved value was almost 2 full integers off in yaw from actual DRMI resonances.  It looks like it was saved when Rana and I were working late last week.  We must have accidentally saved it when it was misaligned, since hysteresis can't do that much.

I want to check the phases for POX and POY with arm locking, just in case.  Also need to set the AS110 phase (which is plugged into the AS11 channels - need to fix the channel names).

 

ELOG V3.1.3-